INAF queue observing, 2018 Jan 28-29

We (the observers) take this space to bring attention of the world to the the awkward situation of large part of research personnel in INAF. We are now performing observations for INAF researchers with LBT, one of the biggest telescope in the world. But we are also valued for our scientific contribution, for results that we publish as first authors and for those in which we are involved.

However, we believe that like for many other INAF post-docs, our contract situation and our expectations do not reflect the great results that we, and INAF thanks to us, are obtaing. Like many other post-docs, we are paid on contracts renewed every year pending the availability of external funds.

Moreover, we are not entitled to officially say a word on institutional choice and in many aspects we are not considered as INAF employees, but like external collaborators. Indeed, we cannot even lead INAF projects and build a independent research.

Now, we believe INAF has given the power and instruments to start changing this situation.

Observer: R. Carini, A. Rossi
Telescope Operator: G. Bechetti
Support Astronomer: O. Kuhn
Instrument: LUCI2

Summary

At the sunset the sky is clear. During the night the seeing get worst. down time: 05:15-05.30 poor seeing for AO; 07:00-13:15 We have taken some date during the down time, who knows if they could be useful.

instruments: LBC works, LUCI1 only imaging and AO, LUCI2 imaging and LS and MOS, MODS1 works, MODS2 only blue channel

Night duration (12\xB0 el twilight) 2:00am- 13:15pm UT = 11 h 15m

Observing time: 2 h 55m (26%)

Weather time loss: 6 h 51m (61%)

technical time loss 1h 30m (13%) (1h AO problems, see AO log )

Data Summary

Project_RUN Objects Instrument Filter ExpTime Notes
darks Antoniucci LUCI2      
darks Vanzella LUCI1      
Vanzella JDz11 LUCI1-AO Ks 1 h  
telluric HIP17971 LUCI2 J, H    
Nisini field3 NGC1333 LUCI2 H 30 m 1 star on top is out of slit
telluric HIP34612 LUCI2 jzspec, HKspec    
Antoniucci IPTF15afq LUCI2 HKspec 13m spectra completed, miss ob ima
Antoniucci O3,O4 RW Aur LUCI2 Hspec, jzspec 0.4+0.1 to be repeated for bad seeing end flexure compensation off
Rossi sn-grb171205A LUCI2 1 h zjspec completed
telluric A3 HIP54608 LUCI2 HKspec    
Piconcelli A3 J1025+2454 LUCI2 Hkspec 28 m 7 spectra with bad seeing
flat and arc Antoniucci LUCI2 zjspec, HKspec, FeII ls + ima

Details

* Darks fro Antoniucci LUCI2: xztau dr tau
  • luci2.20180128.0001-0005 DIT 30 NDIT 1 MER
  • luci2.20180128.0006-0010 DIT 3 NDIT 2 LIR int
  • luci2.20180128.0011-0015 DIT 4 NDIT 3 LIR int
  • luci2.20180128.0016-0020 DIT 10 NDIT 1 LIR int

* Darks fro Antoniucci LUCI2: agl192 iptf15afq
  • luci2.20180128.0021-0025 DIT 30 NDIT 1 MER int
  • luci2.20180128.0026-0030 DIT 5 NDIT 6 LIR int

* Darks fro Antoniucci LUCI2: xztau dr tau
  • luci2.20180128.0031-0035 DIT 25 NDIT 1 LIR int
  • luci2.20180128.0036-0040 DIT 25 NDIT 1 MER norm
  • luci2.20180128.0041-0042 DIT 300 NDIT 1 MER norm

* Darks LUCI1 Vanzella DIT 30 NDIT 12 lir int
  • luci1.20180129.0001-0005 DIT 30 NDIT 2 LIR int

02:30 Nisini program, Field 3. send the script for the telluric
  • acq luci2.20180129.0044-0045
  • the compensation mirror is not aligned, we try the alignment luci2.0046-0047 SA is fixing it

03:00 start of the night, but we are stop for the technical issue.

03:16 the problem with the compensation mirror is not resolved, we slew to LUCI1 - AO, but we have the same problem, we try to fix it.

03:32 we have aligned the field, we start with Vanzella program NCPA off
  • seeing 0.8''-0.9''
  • 03:41 loop is closed, measured magnitude is 13.4
  • sci luci1.20190129.0016-0028
  • luci1.0019-0027 FWHM about 0.2''
  • 04:00 get in pause, AdSec failed
  • 03:07 send the preset again.
  • 03:12 loop closed. We skip to the PSF star offset because it is more than 20 minutes from the beginning of the integration of the script
  • sci luci1.0029-0030
  • 03:17 offset failed.
  • 03:18 we start again sci==luci1.0031-0032==
  • 03:20 failed again the offset
  • 03:22 start again,==luci1.0033== and failed
  • 04:23 send the preset again

03:27 science start again luci1.0034-0056 luci1.0034 maybe is not good
  • seeing 0.8'', measured magnitude 13.2
  • luci1.0038 FWHM =0.2 (upper limit)
  • luci1.0040 fwhm=0.24 (upper limit)
  • luci1.0041 fwhm=0.17 (upper limit)
  • 04:46 seeing 0.7''
  • luci1.0046 fwhm= 0.26 (upper limit)
  • luci1.0048 has sky backgroud too high, maybe is not good
  • luci1.0051 fwhm=0.21''
  • the offset for this field is large, failed the AO at 04:09 and again 04:12
  • 04:15 still waiting for AO
  • 04:19 start again but the AO failed
  • 04:21* there is a problem in the telescope that can not keep the Ref AO star when we integrate the Field B. we have to send the preset again
  • 04:28 loop closed, start the integration luci1.0057
  • we try again the the PSF field luci1.0058-0060
  • fwhm on the pSF star is 0.23'', seeing on the DIMM is 0.75''

04:37 Since in the PSF the loop is suspended often, we pass to the target Field
  • sci =luci1.0061-0062
  • 04:45 loop is fault, the image luci1.0063 is not good
  • 04:57 start again but failed

04:48 send the preset again
  • 04:51 start the integration
  • sci luci1.0064-0066
  • 04:58 pause
  • 05:00 start again
  • sci luci1.0067-0077
  • 05:08 seeing on the DIMM 0.81'', luci1.0072 fwhm 0.24'''
  • 05:14 loop closed, bump of seeing up to 1.5''. start again the integration
  • 05:16 the seeing is getting worst but is stable and about 1.0'', there are thin layers on the horizon
  • luci1.0078-0079
  • 05:20 lost loop, the conditions are too bad. seeing 1.7'''

05:24 we slew to LUCI2, while we were integrating with LUCI1, SA have fixed the problem with LUCI2
  • we start with the telluric of Nisini program field 3
  • 05:36 send the preset
  • acq luci2.0065-0067
  • seeing 1.0''-1.4'' and variable
  • spe luci2.0068-0075

05:51 slew to the field 3
  • acq luci2.0076-0079
  • sci luci2.0080 K band
  • 06:20 seeing variable around 1.4''
  • sci luci2.0081-0085

07:00 slew to the telluric for Antoniucci IPTF15afq
  • acq luci2.0086-0088
  • sci luci2.0089-0092
  • seeing 2''-2.6''

07:17 slew to the target IPTF15afq (we have to finish the HKspec filter )
  • acq luci2.0093-0096
  • 07:27 start sci luci2.0097-0100
  • 07:28 seeing 1.8''-2.0'' high variable

07:46 slew to target RW Aur
  • acq luci2.0101-0104
  • 07:54 start integration, seeing about 3.00''
  • sci luci2.0105-0112 data no good, no flexure compensation during the integration, and the seeing about 3''-4''
  • ima luci2.0113--0116 data no good

08:07 The compensation mirror of LUCI2 is misaligned again

08:15 slew to the standard, but the seeing is too bad and it is difficult the collimation.

08:20 seeing is about 3.0''. SA fix the flexure compensation
  • we align the compensation mirror

08:43 slew to a super back-up program of Rossi, target sn-grb171205A
  • preset failed

08:50 sent the preset again of the telluric
  • acq ==luci2.0121-0123==sngrb170205A
  • sci luci2.0124-0125

09:01 slew to the target sngrb170205A
  • acq luci2.0126-0129
  • sci luci2.0130-0134 windy and seeing > 2.5

09:53 sent again the script
  • sci luci2.0135-0139 seeing 3.4''on the DIMM

10:23 TO is checking the wind

10:45 we can go ahead. Piconcelli A3, standard HIP54608
  • acq luci2.0140-0142
  • sci luci2.0143.-0144
  • seeing variable between 1.2'' and 1.7''

10:57 sent the target A3 script , preset failed, guiding star lost.
  • 11:00 sent again the script.
  • acq luci2.0145-0148
  • sci luci2.0149-0155
  • seeing variable and bad, up to 4.0''
  • 11:35 seeing 1.7'' -2.0''

11:54 slew to Gilli program
  • telluric HIP 56736
  • acq ==luci2.0156-0158
  • sci luci2.0159-0162
  • seeing 1.7''-1.9'', the wind is rising

12:08 Gilli J
  • preset failed, TO checks pointing.
  • 12:11 sent preset again, but failed, the guide star is too faint for this bad seeing.

12:23 In the field of J there is not a bright guide star, we slew to Gilli M
  • acq luci2.0163-==0166
  • seeing is too bad, we can not see the target..seeing 4''-5'', we can not do anything...we slew to the calibrations!!!

*calibration for Antoniucci program
  • flat ima FeII luci2.0167-0177
  • flat HKspec slit 0.75''
  • flat jzspec slit 0.75''
  • arc zjspec (Neon+ Argon)
  • flat zjspec slit 1.0'' luci2.0205-0209

*calibration Nisini Field 3 (flat + arc)
  • ==luci2.0210-0223

*calibration Rossi * flat + arc, slit (1.5''+2.0'')=0224-0237

-- %USERSIG{AndreaRossi - 2018-01-29}%

Comments

 
Topic revision: r11 - 30 Jan 2018, AndreaRossi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback